# Toward Dignity-Aware AI: Next-Generation Elderly Monitoring from Fall Detection to ADL

### v1

@@ 1.1 Motivation
The rapid demographic shift toward aging societies has intensified the urgency of developing intelligent, privacy-preserving monitoring systems to support independent living for older adults. According to the World Health Organization (WHO), falls are one of the leading causes of injury-related mortality among elderly populations, with nearly million falls requiring medical intervention each year [WHO2021Falls]. However, falls are only one dimension of Activities of Daily Living (ADL), which include essential routines such as mobility, sleep, toileting, eating, and personal hygiene [Katz1963ADL]. Continuous monitoring of ADL not only enables timely detection of critical events such as falls, but also provides early indicators of frailty, cognitive decline, or deteriorating health status [ravi2017deep, igarashi2020adl]. This paper advocates for a shift in focus: from isolated fall detection to holistic ADL recognition, supported by edge AI and Federated Learning (FL).
@@ 1.2 Fall Detection as a Proxy Task
Despite the importance of ADL monitoring, most existing studies have focused on fall detection. This is partly because falls are acute and life-threatening events, and partly because benchmark datasets exist. For instance, the SISFall dataset provides annotated accelerometer and gyroscope data for simulated falls [Guerrero2019SISFall]. In this work, we emphasize that fall detection is not our ultimate goal. Instead, we use it as a proxy task to validate algorithms and system components in the absence of large-scale ADL datasets. All preliminary experiments in this paper are conducted solely on SISFall and its GAN-augmented variants. ADL-specific datasets are still under collection in our smart-room testbed.
@@ 1.3 Approaches to Fall Detection
Wearables: Accelerometer and gyroscope sensors embedded in smartphones or body-worn devices are among the most widely studied for fall detection [mubashir2013fallsurvey, Chen2020WearableHAR]. These approaches are generally accurate in controlled conditions. However, they face practical challenges: elderly users may resist wearing sensors, forget to recharge devices, or experience discomfort. Moreover, continuous usage can be intrusive in daily life.
Vision-based: Camera systems, including depth sensors (e.g., Microsoft Kinect), have been applied for fall recognition by analyzing human poses and motion trajectories [Mastorakis2014VisionBased, Stone2015KinectFalls]. Although vision-based systems can achieve high accuracy, they raise serious privacy concerns in sensitive environments such as nursing homes or private homes. In addition, performance may degrade under poor lighting or occlusion, and deployment often requires expensive infrastructure.
Device-free RF: Recent advances in device-free sensing exploit wireless signals, such as WiFi Channel State Information (CSI), Ultra-wideband (UWB), or mmWave radar, to detect falls without requiring wearables or cameras [Wang2017DeviceFree, Adib2015RFTrack, yang2020uwb, zhao2018mfd]. These approaches offer privacy advantages and can penetrate occlusions. However, they often require specialized hardware and are highly sensitive to environmental variations, limiting their scalability.
@@ 1.4 Toward ADL Recognition
While fall detection is a critical starting point, it does not capture the full picture of elderly well-being. ADL monitoring enables a more comprehensive understanding of daily routines, such as whether an elderly person gets out of bed, visits the bathroom, or remains inactive for extended periods [bates2021adlreview, ignatov2018cnn]. Deviations in such patterns may indicate emerging health risks. Thus, our vision is to extend beyond fall detection to a holistic ADL monitoring framework. We stress that our position paper proposes ADL recognition as the ultimate goal, with fall detection serving as an initial case study.
@@ 1.5 Challenges: Data Scarcity and Generalization
A central obstacle is the scarcity of real-world ADL datasets. Data collection in elderly populations raises ethical, privacy, and logistical challenges. While fall datasets such as SISFall exist, large-scale ADL datasets remain rare. To mitigate this, Generative Adversarial Networks (GANs) have been adopted for data augmentation, synthesizing rare falls or subtle transitions to improve class balance and generalization [xu2021fedhar, zhang2019timeseriesgan]. Nevertheless, synthetic data cannot fully replace real-world ADL data, underscoring the need for continued dataset development.
Generalization across heterogeneous environments also poses challenges. Models trained in one home or facility may not transfer well to others due to differences in sensor placement, architecture, or individual behavior. Addressing domain shift requires techniques such as Continual Learning and domain adaptation [delange2021continual, zhuang2020survey].
@@ 1.6 Edge AI and Federated Learning
Privacy concerns further constrain centralized data collection. Federated Learning (FL) provides a paradigm where each device trains locally and shares only model parameters, preserving privacy while enabling collective learning [mcmahan2017communication, Li2020FLChallenges]. Applications of FL in healthcare demonstrate its potential to balance personalization and global generalization [xu2021fedhar, fedhealth2020, rashid2022fedsens]. Moreover, deploying models on embedded devices such as NVIDIA Jetson Orin Nano enables low-latency, energy-efficient inference at the edge [howard2017mobilenets, han2016deeplearningcompression]. In our preliminary experiments, quantized Conv-GRU models achieved efficient inference on Jetson hardware.
@@ 2.1 System Overview
Federated Learning (FL) to collaboratively train global models without centralizing raw data [20,21].
@@ 2.4 Synthetic Data and Data Scarcity
To address class imbalance, we employ conditional GANs (cGANs) trained on SISFall to synthesize rare fall patterns (Figure 4). While the figure depicts the general GAN setup, note that LiDAR-derived inputs mentioned in earlier drafts are not included in current experiments.
@@ 3.3 Model Evaluation Under IID and Non-IID
We implemented and compared three architectures: Conv-GRU+GAT (conceptual extension), CNN-LSTM, and GRU-only. In practice, SISFall IMU streams were used, with GAN-augmented samples included. Evaluation considered two conditions: (i) IID (random partitions) and (ii) Non-IID (cross-partition to simulate inter-site heterogeneity). Table 1 summarizes the results.
@@ 3.5 Embedded Deployment on Jetson Orin Nano
We deployed the Conv-GRU baseline (without multimodal extensions) on a Jetson Orin Nano device. Models were quantized to INT using TensorRT. Fig. 9 shows that inference can be achieved in real time under W power budgets, confirming feasibility for on-device deployment.
@@ 4 Discussion and Future Challenges
Our preliminary results based on SISFall and its GAN-augmented variants confirm that edge-deployed sequence models (e.g., Conv-GRU) can achieve real-time feasibility on embedded devices and maintain robustness under simulated Non-IID and sensor dropout conditions. These findings are promising, but they also highlight the limitations of evaluating on a single proxy dataset. The next stage of our research is to bridge from fall detection on SISFall to holistic ADL recognition in real smart-room deployments. In this section, we outline the key challenges and future directions, linking them directly to our preliminary results.
@@ 4.1 From Proxy Task to ADL Monitoring
While SISFall allows us to test algorithms and deployment feasibility, it does not cover the full spectrum of ADL behaviors (e.g., bed exit/entry, prolonged inactivity, toileting patterns). Thus, our GAN-augmented results should be viewed as methodological validation rather than final benchmarks. Future experiments will require multimodal ADL datasets currently under collection in our smart-room testbed. In this regard, the methodological insights from our figures directly inform future ADL deployments:
@@
By linking these proxy-task results to ADL-oriented challenges, our work goes beyond technical validation and aligns with the broader agenda of supporting independence and dignity in aging societies. This responds directly to expert feedback that emphasizes the necessity of framing fall detection as an intermediate step toward holistic Agetech solutions.
4.2 Sensor Deployment and Maintenance
Our preliminary robustness tests with simulated sensor dropout show that performance degrades under missing inputs. In real deployments, sensors may fail due to noise, occlusion, or hardware degradation. Self-calibration techniques and redundancy in sensor networks will be necessary to sustain performance.
@@ 4.4 Cross-Site Generalization and Domain Shift
Our Non-IID experiments on SISFall partitions highlight the difficulty of cross-site generalization. In practice, differences in room layouts, sensor placements, and individual behaviors will induce significant domain shift. Continual learning and transfer learning strategies will be critical to adapt models across sites and over time [delange2021continual, zhuang2020survey].
@@ 5 Conclusion
In this position paper, we have argued that elderly monitoring research must evolve from isolated fall detection toward holistic ADL recognition. While fall detection remains a critical application, our long-term vision is to monitor daily routines such as sleeping, mobility, and toileting, providing early indicators of frailty and supporting independent living with dignity. Our preliminary results, based solely on the SISFall dataset and its GAN-augmented variants, demonstrate three key aspects: (i) the feasibility of GAN-based data augmentation for balancing rare fall events, (ii) robustness of sequence models under Non-IID conditions, and (iii) real-time inference capability on embedded Jetson Orin Nano devices after quantization. These results validate our methodological approach but must be seen as a proxy-task demonstration rather than final outcomes. No ADL-specific datasets have yet been used, and multimodal sensing (e.g., LiDAR, motion sensors, pressure mats) remains future work.

### v2

@@ 1.1 Motivation
The rapid demographic shift toward aging societies has intensified the urgency of developing intelligent, privacy-preserving monitoring systems to support independent living for older adults. According to the World Health Organization (WHO), falls are one of the leading causes of injury-related mortality among elderly populations, with nearly million falls requiring medical intervention each year [20]. However, falls are only one dimension of Activities of Daily Living (ADL), which include essential routines such as mobility, sleep, toileting, eating, and personal hygiene [10]. Continuous monitoring of ADL not only enables timely detection of critical events such as falls, but also provides early indicators of frailty, cognitive decline, or deteriorating health status [16, 18]. This paper advocates for a shift in focus: from isolated fall detection to holistic ADL recognition, supported by edge AI and Federated Learning (FL).
@@ 1.2 Fall Detection as a Proxy Task
Despite the importance of ADL monitoring, most existing studies have focused on fall detection. This is partly because falls are acute and life-threatening events, and partly because benchmark datasets exist. For instance, the SISFall dataset provides annotated accelerometer and gyroscope data for simulated falls [6]. In this work, we emphasize that fall detection is not our ultimate goal. Instead, we use it as a proxy task to validate algorithms and system components in the absence of large-scale ADL datasets. All preliminary experiments in this paper are conducted solely on SISFall and its GAN-augmented variants. ADL-specific datasets are still under collection in our smart-room testbed.
@@ 1.3 Approaches to Fall Detection
Wearables: Accelerometer and gyroscope sensors embedded in smartphones or body-worn devices are among the most widely studied for fall detection [15, 3]. These approaches are generally accurate in controlled conditions. However, they face practical challenges: elderly users may resist wearing sensors, forget to recharge devices, or experience discomfort. Moreover, continuous usage can be intrusive in daily life.
Vision-based: Camera systems, including depth sensors (e.g., Microsoft Kinect), have been applied for fall recognition by analyzing human poses and motion trajectories [13, 17]. Although vision-based systems can achieve high accuracy, they raise serious privacy concerns in sensitive environments such as nursing homes or private homes. In addition, performance may degrade under poor lighting or occlusion, and deployment often requires expensive infrastructure.
Device-free RF: Recent advances in device-free sensing exploit wireless signals, such as WiFi Channel State Information (CSI), Ultra-wideband (UWB), or mmWave radar, to detect falls without requiring wearables or cameras [19, 1, 21, 23]. These approaches offer privacy advantages and can penetrate occlusions. However, they often require specialized hardware and are highly sensitive to environmental variations, limiting their scalability.
@@ 1.4 Toward ADL Recognition
While fall detection is a critical starting point, it does not capture the full picture of elderly well-being. ADL monitoring enables a more comprehensive understanding of daily routines, such as whether an elderly person gets out of bed, visits the bathroom, or remains inactive for extended periods [2, 9]. Deviations in such patterns may indicate emerging health risks. Thus, our vision is to extend beyond fall detection to a holistic ADL monitoring framework. We stress that our position paper proposes ADL recognition as the ultimate goal, with fall detection serving as an initial case study.
@@ 1.5 Challenges: Data Scarcity and Generalization
A central obstacle is the scarcity of real-world ADL datasets. Data collection in elderly populations raises ethical, privacy, and logistical challenges. While fall datasets such as SISFall exist, large-scale ADL datasets remain rare. To mitigate this, Generative Adversarial Networks (GANs) have been adopted for data augmentation, synthesizing rare falls or subtle transitions to improve class balance and generalization [11, 22]. Nevertheless, synthetic data cannot fully replace real-world ADL data, underscoring the need for continued dataset development.
Generalization across heterogeneous environments also poses challenges. Models trained in one home or facility may not transfer well to others due to differences in sensor placement, architecture, or individual behavior. Addressing domain shift requires techniques such as Continual Learning and domain adaptation [5, 25].
@@ 1.6 Edge AI and Federated Learning
Privacy concerns further constrain centralized data collection. Federated Learning (FL) provides a paradigm where each device trains locally and shares only model parameters, preserving privacy while enabling collective learning [14, 12]. Applications of FL in healthcare demonstrate its potential to balance personalization and global generalization [11, 4, 24]. Moreover, deploying models on embedded devices such as NVIDIA Jetson Orin Nano enables low-latency, energy-efficient inference at the edge [8, 7]. In our preliminary experiments, quantized Conv-GRU models achieved efficient inference on Jetson hardware.
@@ 2.4 Synthetic Data and Data Scarcity
To address class imbalance, we employ conditional GANs (cGANs) trained on SISFall to synthesize rare fall patterns (Figure 4). While the figure depicts the general GAN setup, note that the current experiments focus exclusively on IMU-based SISFall inputs, and multi-modal GAN extensions will be explored in future smart-room studies.
@@ 3.3 Model Evaluation Under IID and Non-IID
We consider several candidate architectures for future ADL-oriented deployments, including Conv-GRU (baseline), GRU-only, CNN-LSTM, and a planned Conv-GRU+GAT extension for multi-sensor fusion. In this position paper, however, all quantitative experiments are conducted on the SISFall IMU proxy task using the Conv-GRU baseline, with GAN-augmented samples included. Evaluation considered two federated conditions: (i) IID (random partitions) and (ii) Non-IID (cross-partition to simulate inter-site heterogeneity). Table 1 summarizes representative results.
@@ 3.5 Embedded Deployment on Jetson Orin Nano
We deployed the Conv-GRU baseline (without multimodal extensions) on a Jetson Orin Nano device. Models were quantized to INT using TensorRT. Fig. 9 shows that real-time inference is achievable for the Conv-GRU baseline within the device’s low-power operating mode. For completeness, projected latency/model-size references for planned extensions (e.g., Conv-GRU+GAT) are also included in the figure, although they are not evaluated in the current SISFall experiments.
@@ 4 Discussion and Future Challenges
Our preliminary results based on SISFall and its GAN-augmented variants confirm that edge-deployed sequence models (e.g., Conv-GRU) can achieve real-time feasibility on embedded devices and maintain robustness under simulated Non-IID conditions, while robustness to sensor dropout remains future work. These findings are promising, but they also highlight the limitations of evaluating on a single proxy dataset. The next stage of our research is to bridge from fall detection on SISFall to holistic ADL recognition in real smart-room deployments. In this section, we outline the key challenges and future directions, linking them directly to our preliminary results.
@@ 4.1 From Proxy Task to ADL Monitoring
While SISFall allows us to test algorithms and deployment feasibility, it does not cover the full spectrum of ADL behaviors (e.g., bed exit/entry, prolonged inactivity, toileting patterns). Thus, our GAN-augmented results should be viewed as methodological validation rather than final benchmarks. Future experiments will require multimodal ADL datasets currently under collection in our smart-room testbed. In this regard, the methodological insights suggested by our figures provide useful guidance for future ADL deployments:
@@
By linking these proxy-task results to ADL-oriented challenges, our work moves beyond technical validation and aligns with the broader agenda of supporting independence and dignity in aging societies. More broadly, this framing reinforces that fall detection should be understood as an intermediate step toward holistic Agetech solutions, rather than an end goal in itself.
@@ 4.3 Cross-Site Generalization and Domain Shift
Our Non-IID experiments on SISFall partitions highlight the difficulty of cross-site generalization. In practice, differences in room layouts, sensor placements, and individual behaviors will induce significant domain shift. Continual learning and transfer learning strategies will be critical to adapt models across sites and over time [5, 25].
@@ 5 Conclusion
In this position paper, we have argued that elderly monitoring research must evolve from isolated fall detection toward holistic ADL recognition. While fall detection remains a critical application, our long-term vision is to monitor daily routines such as sleeping, mobility, and toileting, providing early indicators of frailty and supporting independent living with dignity.
Our preliminary results, based solely on the SISFall dataset and its GAN-augmented variants, demonstrate three key aspects: (i) the feasibility of GAN-based data augmentation for balancing rare fall events, (ii) robustness of sequence models under Non-IID conditions, and (iii) real-time inference capability on embedded Jetson Orin Nano devices after quantization. These results validate our methodological approach but must be seen as a proxy-task demonstration rather than final outcomes. No ADL-specific datasets have yet been used, and multimodal sensing (e.g., LiDAR, motion sensors, pressure mats) remains future work.